### v1

# Semiparametric robust estimation of population location

Abstract
Real‑world measurements often comprise a dominant signal contaminated by a noisy background. Robustly estimating the dominant signal in practice has been a fundamental statistical problem. Classically, mixture models have been used to cluster the heterogeneous population into homogeneous components. Modeling such data with fully parametric models risks bias under misspecification, while fully nonparametric approaches can dissipate power and computational resources. We propose a middle path: a semiparametric method that models only the dominant component parametrically and leaves the background completely nonparametric, yet remains computationally scalable and statistically robust. So instead of outlier downweighting, traditionally done in robust statistics literature, we maximize the observed likelihood such that the noisy background is absorbed by the nonparametric component. Computationally, we propose a new approximate FFT-accelerated likelihood maximization algorithm. Empirically, this FFT plug‑in achieves order‑of‑magnitude speedups over vanilla weighted EM while preserving statistical accuracy and large sample properties.
1 Introduction
Finite mixture models have been a standard tool for modeling heterogeneous populations, discovering subpopulations, and performing model‑based clustering. Classical formulations impose a common parametric family on all components (e.g., Gaussian mixtures) and rely on identifiability assumptions or parsimonious parameterizations to render the estimation problem well‑posed [9, 4, 1, 8]. The EM algorithm is the workhorse in practice [32, 24, 10, 13]. Yet the assumptions that make parametric mixtures tractable—common family, structured covariances, or a fixed number of components—are almost never aligned with actual noisy datasets. In many applications the investigator primarily cares about the most dominant signal peak (its location and mass), while the remainder of the distribution is a complicated amalgam of weaker modes, tails, or background phenomena that defy clear parametric specification. For such tasks, fitting a fully parametric KK‑component model can be brittle: identifiability hinges on modeling all components correctly, and estimation can deteriorate due to over‑fitting or weak identifiability [31, 7].
A parallel line of work relaxes parametric assumptions on one component. In two‑component mixtures where one distribution is known and the other is nonparametric (sometimes with symmetry or shape constraints), there are strong identifiability and estimation results for the mixing proportion and for features of the unknown component; see, e.g., [4] and [25]. In machine learning, the related mixture‑proportion estimation problem has been studied in semi‑supervised novelty detection and positive‑unlabeled learning, with estimators and rates under separability or irreducibility conditions [3, 27, 26]. These strands underscore a theme: if the inferential target is a weight or a single component, strict distributional modeling of the remainder can be avoided.
Our paper embraces this philosophy and pushes it in a direction that is distinct from classical robust statistics. Robust methods à la Huber and Hampel downweight or reweight observations to resist the effect of outliers—with M‑estimators, influence‑function design, or divergence‑based criteria—but still typically posit a parametric data‑generating model and optimize a single loss over all data. Robust clustering extensions either replace Gaussian errors by heavier‑tailed $t$ components or trim a fraction of points to curb spurious solutions. While powerful for outlier resistance, these approaches differ in motivation and technique from our objective of extracting the dominant peak and its mass without modeling the rest [18, 16, 2].
We consider a semiparametric “contamination” mixture $f(y)\;=\;\pi_{0}\,f_{0}(y;\theta)\;+\;(1-\pi_{0})\,f_{\mathrm{bg}}(y),\qquad\pi_{0}\in(0,1),$ in which $f_{0}(\cdot;\theta)$ is a parametric family for a dominant component (e.g., a location or location–scale family) and the background density $f_{\mathrm{bg}}^{*}$ is left completely unrestricted. The inferential target is the dominant component’s location parameter $\theta$. We do not attempt full identifiability of a multi‑component parametric mixture; rather, our goal is targeted recovery of the dominant component under certain regularity. This formulation sits at the intersection of semiparametric mixtures “with one unknown component” and mixture‑proportion estimation [4, 3, 27, 25, 26].
Algorithmically, we maximize a plug–in log‑likelihood where the unknown background $f_{\mathrm{bg}}^{*}$ is replaced by a kernel density estimator (KDE) computed once per EM iteration via an FFT‑accelerated convolution on a uniform grid with linear binning and zero padding [28, 29, 35, 36]. This reduces the nonparametric update from $O(nG)$ to $O(G\log G)$ for $G$ grid points, while off‑grid evaluations use linear interpolation whose error we bound by the usual $C^{2}$ truncation rate $O(\Delta^{2}/h^{3})$ (here $\Delta$ is the grid spacing and $h$ the bandwidth), so the FFT/binned criterion differs from the exact plug–in criterion by a uniform $o_{p}(1)$ term.
On the statistical side, we establish (i) consistency of the FFT plug–in estimator $\hat{\theta}_{n}$ under (a) standard smoothness/undersmoothing for KDE, (b) a windowed body–tail control for the log‑likelihood, and (c) a local curvature (strong concavity) condition for the population objective around $\theta^{*}$; and (ii) a CLT for smooth functionals after strengthening the discretization to $\sqrt{n}\,\varepsilon_{n}\!\to\!0$, where $\varepsilon_{n}\!=\!\tfrac{\|K^{\prime\prime}\|_{\infty}}{4}\Delta_{n}^{2}/h_{n}^{3}$ is the binned‑FFT interpolation error. The proofs combine a uniform oracle LLN for the parametric part [34] with (a) uniform control of the FFT/KDE approximation on a growing body window and (b) crude but sufficient tail envelopes that shrink against $\log(1/h_{n})$. The resulting consistency and CLT coincide with “what one would expect” if the background density were known, modulo the small plug–in remainder.
Classical robust statistics mitigates sensitivity by modifying the loss, bounding the influence function, or minimizing a robust divergence [19, 15, 23]. Our approach differs in these aspects: (1) we do not posit a global parametric model for all observations—only the dominant component is parametric, the rest is nonparametric; (2) the weights arise from posterior membership in a mixture rather than from a downweighting scheme engineered via an influence curve. In this sense our method is closer to the semiparametric mixture literature, except here the unknown component is neither known nor shape‑constrained, and we leverage a fast nonparametric plug‑in evaluated inside the EM algorithm.
2 Estimating a single dominant component
We first consider the problem of estimating a single dominating subpopulation from a heterogeneous population. Towards this, we model the data $Y_{1},\dots,Y_{n}$ through a two component mixture $f,$ which can be written as
$f(y)\;=\;\pi_{0}f_{0}(y;\theta)+(1-\pi_{0})f_{1}(y),$ (1)
where $f_{0}(\cdot\,;\theta)$ is our main component of interest and it is modeled through a parametric family with unknown parameter $\theta\in\Theta\subset\mathbb{R}^{q},$ and $f_{1}$ is kept unspecified, modeling the remaining mixture of the other background components. The mixing weights $\pi_{0}$ and $1-\pi_{0}$ are unknown but constant. We observe $n$ i.i.d. samples $\mathbf{y}=(y_{1},\cdots,y_{n})$ from this mixture.
The average log-likelihood function of (1) is
$\mathcal{L}(\Theta;\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}\log\Bigl\{\pi_{0}f_{0}(y_{i};\theta)+(1-\pi_{0})f_{1}(y_{i})\Bigr\},\qquad\Theta=(\pi_{0},\theta)$ (2)
We cannot optimize the above expression directly because $f_{1}(y)$ is unknown. For this reason we approximate $f_{1}$ via a kernel density estimator
$\tilde{f}^{\mathrm{exact}}_{1,n}(y;\omega)=\frac{1}{h_{n}}\sum_{j=1}^{n}\omega_{j}\,\mathbb{K}\!\left(\frac{y_{j}-y}{h_{n}}\right),$ (3)
where $\mathbb{K}(\cdot)$ is a kernel function satisfying Condition 1, $\omega_{j}$’s are weighting parameters such that $\sum_{i=1}^{n}\omega_{j}=1$ and $h_{n}>0$ is the bandwidth. $\omega_{j}$ is a function of the observed data $y_{j}.$ For notational simplicity we drop this explicit dependence in our expressions.
Condition 1 (Kernel properties).
The kernel function $\mathbb{K}(z)$ satisfies
$\int\mathbb{K}(z)\,dz=1,\quad\int\mathbb{K}(z)\,z\,dz=0,\quad\int\mathbb{K}(z)\,z^{2}\,dz<\infty.$
The working log-likelihood using this plug-in estimate for the nonparametric density $\tilde{f}_{1}$ is
$\tilde{\mathcal{L}}_{n}(\Theta,\omega;\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}\log\Big\{\pi_{0}f_{0}(y_{i};\theta)+(1-\pi_{0})\tilde{f}^{\mathrm{exact}}_{1,n}(y_{i})\Big\},\qquad\Theta=(\pi_{0},\theta).$
Let $\Theta^{*}$ denote the truth of $\Theta$, and $f_{1}^{*}$ denote the truth of $f_{1}$ then from [37] we have the following theorem
Theorem 1.
[…]
Under standard regularity and boundedness conditions  [37] guarantee the identifiability of $\hat{\Theta}$ and one can only identify the component that is large enough. The component with very small weight cannot be identified. A concavity assumption further guarantees the uniqueness of the parameter estimate.
The algorithm for the likelihood maximization proposed in [37] requires evaluating $\tilde{f}^{\mathrm{exact}}_{1,n}(y;\omega)=\frac{1}{h_{n}}\sum_{j=1}^{n}\omega_{j}\,\mathbb{K}\!\left(\frac{y_{j}-y}{h_{n}}\right)$ at all $Y_{i}$. This leads to a computational complexity of $\mathcal{O}(n^{2})$ per iteration. In finite samples their algorithm can also suffer from role flips where the parametric block absorbs the tight minority while the KDE models the diffuse majority. To counter such issues, we propose the following ideas.
2.1 Fast weighted KDE via binning and FFT
The kernel density estimator of the unknown component $f_{1}$ is
$\tilde{f}_{1}(y)=\sum_{j=1}^{n}\omega_{j}\,\mathbb{K}_{h_{n}}(y-Y_{j}),\qquad\mathbb{K}_{h_{n}}(u)=\frac{1}{h_{n}}\mathbb{K}\!\left(\frac{u}{h_{n}}\right).$
where $\sum_{i}\omega_{i}=1.$ A direct evaluation of this weighted KDE at all $n$ data points costs $O(n^{2})$ per EM iteration. We exploit that a binned KDE is a discrete convolution and can be computed in $O(M\log M)$ time on a grid through Fast Fourier Transform (FFT).
Linear binning.
Fix an equally spaced grid $x_{1}<\cdots<x_{M}$ with spacing $\Delta$ such that $x_{m}=x_{1}+(m-1)\Delta$, $m=1,\dots,M$ and an array $c[1:M]$ initialized to zeros. For each $Y_{i}$, define $p_{i}=(Y_{i}-x_{1})/\Delta$, $j_{i}=\lfloor p_{i}\rfloor$, $\alpha_{i}=p_{i}-j_{i}\in[0,1)$. We split the weights linearly as:
$c[j_{i}]\!+\!=\omega_{i}(1-\alpha_{i}),\qquad c[j_{i}+1]\!+\!=\omega_{i}\alpha_{i}.$
Equivalently,
[…]
Then $c[m]\geq 0$ and $\sum_{m=0}^{M-1}c[m]=1$.
Discrete convolution on the grid via FFT.
At grid points $x_{m},$ we have $x_{m}=x_{0}+m\Delta$ for $m=0,\ldots,M-1$. Define the sampled kernel $k_{h}[r]:=K_{h_{n}}(r\Delta)$ for integer lags $r\in\mathbb{Z}$. The binned KDE on the grid is the linear convolution
$\tilde{f}^{\mathrm{FFT}}_{1,n}(x_{m})=\sum_{\ell=0}^{M-1}c[\ell]\;k_{h}[m-\ell],\qquad m=0,\ldots,M-1.$
To compute this with an FFT, choose a kernel radius $R\in\mathbb{N}$, set $L:=2R+1$, and pick an FFT length $P\ \geq\ M+L-1\quad\text{(in practice, }P=2^{\lceil\log_{2}(M+L-1)\rceil}\text{).}$ Construct two length‑$P$ arrays:
$\tilde{c}[p]=\begin{cases}c[p],&0\leq p\leq M-1,\\ 0,&M\leq p\leq P-1,\end{cases}\qquad\tilde{k}[p]=\begin{cases}k_{h}[0],&p=0,\\ k_{h}[p],&1\leq p\leq R,\\ k_{h}[-(P-p)],&P-R\leq p\leq P-1,\\ 0,&\text{otherwise}.\end{cases}$
The negative lags $k_{h}[-1],\ldots,k_{h}[-R]$ are stored at positions $P\!-\!1,\ldots,P\!-\!R$. If $K$ is symmetric then $k_{h}[-r]=k_{h}[r]$.
Compute the PP‑point FFTs and inverse FFT:
$C=\mathrm{FFT}_{P}(\tilde{c}),\qquad K=\mathrm{FFT}_{P}(\tilde{k}),\qquad Z=C\odot K,\qquad s=\mathrm{IFFT}_{P}(Z),$
where $\odot$ denotes elementwise multiplication. With the above padding and placement, the first $M$ entries of $s$ equal the desired linear convolution values,
[…]
Off-grid evaluation.
For sample points $y_{i}$, linearly interpolate between the nearest grid values $\tilde{f}(x_{m})$ values. Define
$\tilde{f}^{\mathrm{FFT}}_{1,n}(y_{i}):=(1-\alpha_{i})\tilde{f}^{\mathrm{FFT}}_{1,n}(x_{j(i)})+\alpha_{i}\tilde{f}^{\mathrm{FFT}}_{1,n}(x_{j(i)+1}).$
Per–iteration cost.
One EM iteration has three steps: (i) linear binning of $n$ weighted points into $M$ grid bins ($O(n)$); (ii) one zero–padded FFT to compute the grid convolution ($O(P\log P)$); (iii) linear interpolation back to the $n$ data points ($O(n)$). The FFT length only needs to satisfy $P\geq M+L-1$, where $L$ is the number of relevant kernel taps on the grid. Because the kernel is localized, $L\ll M$, so $P=M+O(L)=\Theta(M)$ and the FFT stage is $O(M\log M)$. Hence the per–iteration time is
$T_{\mathrm{iter}}=O(n)+O(M\log M)+O(n)=O\!\big(n+M\log M\big),$
with $O(M)$ memory. The grid size scales as $M\approx\mathrm{range}/\Delta_{n}$, so $M$ grows with $n$ only through the chosen grid step $\Delta_{n}$. In practical choices $M=o(n)$, making the overall iteration effectively linear in $n$.
Approximation error bound.
In this FFT implementation due to first binning the data onto a grid, convolving these binned counts with the rescaled kernel $K_{h_{n}}$ on the grid via FFT, and recovering $\tilde{f}^{\mathrm{FFT}}_{1,n}(y)$ for arbitrary $y$ by linear interpolation, we encounter approximation error in two stages. A second–order Taylor expansion shows that both the binning step and the interpolation step introduce at most a $O(\Delta_{n}^{2}/h_{n}^{3})$ error per kernel term. Summing over the weights and using $\sum_{i}\hat{\omega}_{i}=1$ leads to a uniform bound on the discrepancy between the exact weighted KDE $\tilde{f}^{\mathrm{exact}}_{1,n}$ and its FFT approximation $\tilde{f}^{\mathrm{FFT}}_{1,n}$ over both the sample points $Y_{i}$ and the whole grid interval.
Lemma 1.
Assume $\mathbb{K}\in C^{2}(\mathbb{R})$ be bounded with $\|\mathbb{K}^{\prime\prime}\|_{\infty}<\infty$ and $\mathbb{K}_{h_{n}}(u):=h_{n}^{-1}\mathbb{K}(u/h_{n})$. Then
$\max_{1\leq i\leq n}\big|\tilde{f}^{\mathrm{FFT}}_{1,n}(Y_{i})-\tilde{f}^{\mathrm{exact}}_{1,n}(Y_{i})\big|\ \leq\ \frac{1}{4}\,\|K^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
Moreover, from definition of $\tilde{f}^{\mathrm{FFT}}_{1,n}(y)$ through linear interpolation between neighboring grid values, we have
$\sup_{y\in[x_{0},x_{M-1}]}\big|\tilde{f}^{\mathrm{FFT}}_{1,n}(y)-\tilde{f}^{\mathrm{exact}}_{1,n}(y)\big|\ \leq\ \frac{1}{4}\,\|K^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
2.2 Consistency and large sample properties
As stated before, we observe i.i.d. data $Y_{1},\dots,Y_{n}$ from the one–dimensional mixture
$f(y)\;=\;\pi_{0}f_{0}(y;\theta)+\pi_{1}f_{1}(y),\qquad\pi_{0}\in(0,1),\ \pi_{1}:=1-\pi_{0},$
where $\theta^{*}\in\mathcal{A}\subset\mathbb{R}$ and we maximize over a compact $\mathcal{C}\subset\mathcal{A}$ containing $\theta^{*}$.
Theorem 2.
Under Assumptions 1, 2, 3, 4, the location estimate obtained from maximizing the FFT plug-in log likelihood
$\hat{\theta}_{n}\in\arg\max_{\theta\in\mathcal{C}}\tilde{L}_{n}(\theta).$
converges to the true location parameter $\theta^{*}$ in probability, i.e.,
$\hat{\theta}_{n}\ \xrightarrow{p}\ \theta^{*}.$
Theorem 3.
Assume the conditions of Theorem 2 together with Assumptions 6, 7, 8 and 9. Then
$\sqrt{n}\,(\hat{\theta}_{n}-\theta^{*})\ \rightsquigarrow\ \mathcal{N}\!\big(0,\,I(\theta^{*})^{-1}\big).$
2.3 Estimation algorithm
In order to obtain the FFT plug–in estimator $\hat{\theta}$ and the associated weights $\hat{\omega}=(\hat{\omega}_{1},\ldots,\hat{\omega}_{n})^{\top}$ for the nonparametric component, we use the EM algorithm on this approximate FFT plug-in log-likelihood.
Proposition 1.
At iteration $t$, given $(\theta^{(t)},\pi_{0}^{(t)})$ and a current background approximation $\tilde{f}^{\mathrm{FFT},(t)}_{1,n}$, define $\pi_{1}^{(t)}:=1-\pi_{0}^{(t)}$ and perform:
E–step (posterior probabilities).
$r_{i}^{(t)}\ :=\ \frac{\pi_{0}^{(t)}\,f_{0}(Y_{i};\theta^{(t)})}{\pi_{0}^{(t)}\,f_{0}(Y_{i};\theta^{(t)})+\pi_{1}^{(t)}\,\tilde{f}^{\mathrm{FFT},(t)}_{1,n}(Y_{i})},\qquad s_{i}^{(t)}:=1-r_{i}^{(t)},$
where the off–grid evaluation is
$\tilde{f}^{\mathrm{FFT},(t)}_{1,n}(Y_{i})=(1-\alpha_{i})\,\tilde{f}^{\mathrm{FFT},(t)}_{1,n}(x_{j(i)})+\alpha_{i}\,\tilde{f}^{\mathrm{FFT},(t)}_{1,n}(x_{j(i)+1}).$
M–step for $\pi_{0}$ and $\theta$. Let
$Q^{(t)}(\pi_{0},\theta)\ :=\ \frac{1}{n}\sum_{i=1}^{n}\Big\{r_{i}^{(t)}\,[\log\pi_{0}+\log f_{0}(Y_{i};\theta)]+s_{i}^{(t)}\,\log(1-\pi_{0})\Big\}.$
Then
$\pi_{0}^{(t+1)}=\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(t)},\qquad\theta^{(t+1)}\in\arg\max_{\theta}\ \frac{1}{n}\sum_{i=1}^{n}r_{i}^{(t)}\,\log f_{0}(Y_{i};\theta).$
Equivalently, $\theta^{(t+1)}$ solves the weighted score equation
$\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(t)}\,\partial_{\theta}\log f_{0}(Y_{i};\theta)=0.$
Background update (for the next E–step). Normalize the posterior probabilities
$\omega_{i}^{(t)}:=\frac{s_{i}^{(t)}}{\sum_{j=1}^{n}s_{j}^{(t)}},\qquad\omega_{i}^{(t)}\geq 0,\ \sum_{i}\omega_{i}^{(t)}=1,$
bin them to the grid
$c^{(t)}[m]\ :=\ \sum_{i=1}^{n}\omega_{i}^{(t)}\Big[(1-\alpha_{i})\mathbf{1}\{m=j(i)\}+\alpha_{i}\mathbf{1}\{m=j(i)+1\}\Big],\quad m=0,\ldots,M-1,\quad\sum_{m}c^{(t)}[m]=1,$
and compute the grid convolution by zero–padded FFT (linear convolution)
$\tilde{f}^{\mathrm{FFT},(t+1)}_{1,n}(x_{m})=\sum_{\ell=0}^{M-1}c^{(t)}[\ell]\;K_{h_{n}}(x_{m}-x_{\ell}),\qquad m=0,\ldots,M-1.$
For off–grid points $Y_{i}$ in the next E–step use the linear interpolation
$\tilde{f}^{\mathrm{FFT},(t+1)}_{1,n}(Y_{i})=(1-\alpha_{i})\,\tilde{f}^{\mathrm{FFT},(t+1)}_{1,n}(x_{j(i)})+\alpha_{i}\,\tilde{f}^{\mathrm{FFT},(t+1)}_{1,n}(x_{j(i)+1}).$
Our FFT-accelerated approximate EM algorithm for locating the dominant subpopulation location is presented below.
Algorithm 1 FFT accelerated EM for a two–component semiparametric mixture
Data $\{Y_{i}\}_{i=1}^{n}$; kernel $\mathbb{K}$; bandwidth $h$; FFT grid $\{x_{j}\}_{j=0}^{M-1}$; tolerance tol; maximum iterations maxit. Randomly split the data by assigning $r_{i}^{(0)}\sim Ber(0.5).$ Initialize $\pi_{0}^{(0)},\,\theta^{(0)}$ from the M-step in Proposition 1
repeat
  Background update: Normalise weights $\omega_{i}^{(t)}\leftarrow s_{i}^{(t)}\big/\sum_{k=1}^{n}s_{k}^{(t)}$; bin onto $\{x_{j}\}$ and convolve with $K_{h}$ via FFT to get $\{\tilde{f}_{1}^{(t)}(x_{j})\}_{j=0}^{M-1}$
  Off–grid interpolation: For each $i$, find $j(i)$ with $Y_{i}\in[x_{j(i)},x_{j(i)+1})$. $\alpha_{i}=(Y_{i}-x_{j(i)})/\Delta$ and set
$\tilde{f}_{1}^{(t)}(Y_{i})\leftarrow(1-\alpha_{i})\tilde{f}_{1}^{(t)}(x_{j(i)})+\alpha_{i}\tilde{f}_{1}^{(t)}(x_{j(i)+1})$
  E–step (posterior probs):
$r_{i}^{(t)}\leftarrow\frac{\pi_{0}^{(t)}f_{0}(Y_{i};\theta^{(t)})}{\pi_{0}^{(t)}f_{0}(Y_{i};\theta^{(t)})+(1-\pi_{0}^{(t)})\,\tilde{f}_{1}^{(t)}(Y_{i})},\qquad s_{i}^{(t)}\leftarrow 1-r_{i}^{(t)}\quad(i=1,\dots,n)$
  M–step:
$\pi_{0}^{(t+1)}\leftarrow\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(t)},\qquad\theta^{(t+1)}\in\arg\max_{\theta\in\mathcal{C}}\sum_{i=1}^{n}r_{i}^{(t)}\log f_{0}(Y_{i};\theta)$
  Working log–likelihood:
$\tilde{\mathcal{L}}^{(t+1)}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\log\!\big\{\pi_{0}^{(t+1)}f_{0}(Y_{i};\theta^{(t+1)})+(1-\pi_{0}^{(t+1)})\,\tilde{f}_{1}^{(t)}(Y_{i})\big\}$
  $t\leftarrow t+1$
until $|L^{(t)}-L^{(t-1)}|<\texttt{tol}$ or $t=\texttt{maxit}$
Output: $\hat{\theta}\leftarrow\theta^{(t)}$,   $\hat{\pi}_{0}\leftarrow\pi_{0}^{(t)}$.
Notes on practical usage.
To prevent the nonparametric component from latching onto a diffuse majority or a sharp minority, we recommend the following heuristic. First, apply a simple one–dimensional kk–means clustering to the data and split it into “left” and “right” clusters. Then initialize the posterior probabilities $r_{i}^{(0)}$ twice: once using the proportion in the smaller cluster and once using the proportion in the larger cluster. Run Algorithm 1 under both initializations and retain the estimate $\hat{\theta}$ corresponding to the fit with the larger estimated mixing weight $\hat{\pi}_{0}$ on the parametric component.
3 Estimating more than 2 components
[…]
We consider our data $Y_{1},\dots,Y_{n}$ coming from a density $f$ on $\mathbb{R}$ of the form
$f(y)\;=\;\sum_{j=1}^{J}\alpha_{j}f_{0}(y;\theta_{j})\;+\;\Big(1-\sum_{j=1}^{J}\alpha_{j}\Big)\,f_{\mathrm{bg}}(y),\qquad y\in\mathbb{R},$
with $\alpha_{j}\in(0,1)$, $\sum_{j=1}^{J}\alpha_{j}<1$, and $\theta_{j}\in\mathcal{A}\subset\mathbb{R}$. Assume that for some $\ell\geq 2$ we have already obtained consistent estimators $(\hat{\theta}_{j},\hat{\alpha}_{j})$ of $(\theta_{j},\alpha_{j})$ for $j<\ell$.
Residual mixture at stage $\ell$.
[…]
$f^{(\ell)}(y):=\frac{f(y)-\sum_{j<\ell}\alpha_{j}f_{0}(y;\theta_{j})}{1-\sum_{j<\ell}\alpha_{j}},\qquad y\in\mathbb{R}.$
Then $f^{(\ell)}$ is a probability density. $f^{(\ell)}$ admits a further two–component decomposition
$f^{(\ell)}(y)=\alpha_{\ell}^{(\ell)}f_{0}(y;\theta_{\ell})+(1-\alpha_{\ell}^{(\ell)})f_{\mathrm{bg}}^{(\ell)}(y),$ (4)
where
$\alpha_{\ell}^{(\ell)}:=\frac{\alpha_{\ell}}{1-\sum_{j<\ell}\alpha_{j}}\in(0,1),$
and $f_{\mathrm{bg}}^{(\ell)}$ is a (completely unspecified) background density at stage $\ell$. Our target parameter at stage $\ell$ is
$\eta_{\ell}:=(\theta_{\ell},\alpha_{\ell}^{(\ell)})\in\mathcal{C}\times\Pi_{\min},$
where $\mathcal{C}\subset\mathcal{A}$ is compact with $\theta_{\ell}\in\mathcal{C}$ and $\Pi_{\min}:=[\underline{\alpha}_{\ell},1-\underline{\alpha}_{\ell}]$ for some $0<\underline{\alpha}_{\ell}<\alpha_{\ell}^{(\ell)}<1$.
Estimated residual weights and plug–in criterion.
Let $(\hat{\theta}_{j},\hat{\alpha}_{j})$, $j<\ell$, denote the stage–(1,,1)($\dots,\ell-1$) estimators, and let $\hat{g}$ be an estimator of $g$ (for instance, the fitted mixture density from earlier stages). Define the unnormalised residual weights
$\widehat{v}_{i}^{(\ell)}:=1-\sum_{j<\ell}\hat{\alpha}_{j}\,\frac{f_{0}(Y_{i};\hat{\theta}_{j})}{\hat{g}(Y_{i})},\qquad i=1,\dots,n,$
and assume $\widehat{v}_{i}^{(\ell)}\geq 0$ a.s. for all $i$ and large $n$. Normalise to obtain probability weights
$\widehat{w}_{i}^{(\ell)}:=\frac{\widehat{v}_{i}^{(\ell)}}{\sum_{k=1}^{n}\widehat{v}_{k}^{(\ell)}},\qquad i=1,\dots,n.$
These weights approximate sampling from $f^{(\ell)}$: they downweight observations explained by earlier peaks.
[…]
$\ell^{(\ell)}(y;\eta,u):=\log\{\alpha f_{0}(y;\theta)+(1-\alpha)u(y)\},$
[…]
$\tilde{\mathcal{L}}_{n}^{(\ell)}(\eta):=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}\,\ell^{(\ell)}\big(Y_{i};\eta,\tilde{f}_{\ell,n}^{\mathrm{FFT}}\big).$
[…]
$\hat{\eta}_{\ell,n}\in\arg\max_{\eta\in\mathcal{C}\times\Pi_{\min}}\tilde{L}_{n}^{(\ell)}(\eta).$
The corresponding oracle criterion replaces $\tilde{f}_{\ell,n}^{\mathrm{FFT}}$ by $f_{\mathrm{bg}}^{(\ell)}$ and $P_{n}^{(\ell)}:=\sum_{i}\widehat{w}_{i}^{(\ell)}\delta_{Y_{i}}$ by $f^{(\ell)}$:
$\mathcal{L}^{(\ell)}(\eta):=\mathbb{E}_{f^{(\ell)}}\big[\ell^{(\ell)}(Y;\eta,f_{\mathrm{bg}}^{(\ell)})\big],\qquad\eta\in\mathcal{C}\times\Pi_{\min}.$
[…]
$\hat{\eta}_{\ell,n}=(\hat{\theta}_{\ell,n},\hat{\alpha}_{\ell,n})\ \xrightarrow{p}\ \eta_{\ell}^{*}=(\theta_{\ell}^{*},\alpha_{\ell}^{(\ell)}).$
[…]
Suppose the conditions of Theorem 4 hold, together with Assumptions 1517. Then
$\sqrt{n}\big(\hat{\eta}_{\ell,n}-\eta_{\ell}^{*}\big)\ \rightsquigarrow\ \mathcal{N}\big(0,J^{(\ell)}(\eta_{\ell}^{*})^{-1}\big).$
In particular, the marginal distribution of $\sqrt{n}(\hat{\theta}_{\ell,n}-\theta_{\ell}^{*})$ is normal with covariance given by the $(\theta,\theta)$–block of $J^{(\ell)}(\eta_{\ell}^{*})^{-1}$.
[…]
1:Data $\{Y_{i}\}_{i=1}^{n}$; stage $\ell\geq 1$; previous components $(\hat{\alpha}_{j},\hat{\theta}_{j})_{j<\ell}$ (empty if $\ell=1$); kernel $K$ with bandwidth $h_{n}$; FFT grid $\{x_{j}\}_{j=0}^{M-1}$ (spacing $\Delta$); tolerance tol; maximum iterations maxit.
[…]
3:Residual weights
[…]
5:Raw residual scores $\hat{v}_{i}^{(\ell)}\leftarrow\max\{0,\,1-\hat{f}_{1:(\ell-1)}(Y_{i})/\max(\hat{f}_{1:(\ell-1)}(Y_{i}),\varepsilon)\}$.
6:Normalised residual weights $\hat{w}_{i}^{(\ell)}\leftarrow\hat{v}_{i}^{(\ell)}/\sum_{k}\hat{v}_{k}^{(\ell)}$.
7:
8:Initialisation
9:Choose $\theta_{\ell}^{(0)}$ (e.g. robust centre of $\{Y_{i}\}$ w.r.t. $\hat{w}_{i}^{(\ell)}$) and $\alpha_{\ell}^{(\ell,0)}\in(0,1)$.
10:Set $r_{i}^{(\ell,0)}\leftarrow s_{i}^{(\ell,0)}\leftarrow\tfrac{1}{2}$, $t\leftarrow 0$, $\tilde{\mathcal{L}}_{\ell}^{(0)}\leftarrow$-$\infty$.
11:repeat
12:  Background FFT KDE (stage $\ell$)
13:  $\omega_{i}^{(\ell,t)}\leftarrow\hat{w}_{i}^{(\ell)}s_{i}^{(\ell,t)}\Big/\sum_{k}\hat{w}_{k}^{(\ell)}s_{k}^{(\ell,t)}$.
14:  Bin $\{(Y_{i},\omega_{i}^{(\ell,t)})\}$ on $\{x_{j}\}$ and convolve with $K_{h_{n}}$ via FFT to get $\{\tilde{f}_{\ell}^{(\mathrm{FFT},t)}(x_{j})\}$.
15:  Interpolate linearly to obtain $\tilde{f}_{\ell}^{(\mathrm{FFT},t)}(Y_{i})$ for all $i$.
16:
17:  E–step (stage $\ell$)
18:  for $i=1,\dots,n$ do
19:   $f_{i}^{(\ell,t)}\leftarrow\alpha_{\ell}^{(\ell,t)}f_{0}(Y_{i};\theta_{\ell}^{(t)})+(1-\alpha_{\ell}^{(\ell,t)})\tilde{f}_{\ell}^{(\mathrm{FFT},t)}(Y_{i})$.
20:   $r_{i}^{(\ell,t+1)}\leftarrow\alpha_{\ell}^{(\ell,t)}f_{0}(Y_{i};\theta_{\ell}^{(t)})/f_{i}^{(\ell,t)}$,  $s_{i}^{(\ell,t+1)}\leftarrow 1-r_{i}^{(\ell,t+1)}$.
21:  end for
22:
23:  M–step (stage $\ell$)
24:  $\alpha_{\ell}^{(\ell,t+1)}\leftarrow\sum_{i=1}^{n}\hat{w}_{i}^{(\ell)}r_{i}^{(\ell,t+1)}$.
25:  $\theta_{\ell}^{(t+1)}\in\arg\max_{\theta\in\overline{\mathcal{A}}}\sum_{i=1}^{n}\hat{w}_{i}^{(\ell)}r_{i}^{(\ell,t+1)}\log f_{0}(Y_{i};\theta)$.
26:
27:  Working plug–in log–likelihood
28:  $\tilde{\mathcal{L}}_{\ell}^{(t+1)}\leftarrow\sum_{i=1}^{n}\hat{w}_{i}^{(\ell)}\log\!\big\{\alpha_{\ell}^{(\ell,t+1)}f_{0}(Y_{i};\theta_{\ell}^{(t+1)})+(1-\alpha_{\ell}^{(\ell,t+1)})\tilde{f}_{\ell}^{(\mathrm{FFT},t)}(Y_{i})\big\}$.
29:  $t\leftarrow t+1$.
30:until $|\tilde{\mathcal{L}}_{\ell}^{(t)}-\tilde{\mathcal{L}}_{\ell}^{(t-1)}|<\texttt{tol}$ or $t=\texttt{maxit}$
31:Output (stage $\ell$)
[…]
4 Conclusion
This work set out as a robust location estimation problem through likelihood–based mixture estimation, in a different spirit from classical techniques in robust statistics. We formalized a semiparametric contamination model, developed an FFT–accelerated plug–in EM algorithm that estimates the parametric peak while computing the unknown nonparametric background by a fast kernel density estimator. We have also proved consistency and asymptotic normality of the resulting estimates, both in a single–stage setting and in a sequential “peel–off” scheme that extracts multiple dominant modes one by one. Our estimates provide similar statistical large sample guarantees as the usual maximum likelihood estimates. Compared to vanilla EM, our proposed method leverages FFT–plug-in KDE to achieve near–linear or $(G\log G)$ complexity per iteration, making it practical for large-scale problems. Together, these features provide a principled and interpretable way to decompose complex and heterogeneous multi–modal distributions into a small number of well–identified peaks plus an unrestricted remainder, without sacrificing the familiar EM workflow that practitioners find easy to implement and extend.
5 Experiments
We now illustrate the empirical performance of our method on a range of synthetic and real-world datasets. The synthetic examples aim to probe the ability of the algorithm to recover the dominant peak location under varying levels of contamination and multimodality, while keeping full control over the ground truth. We then turn to real datasets where the experimenter has no control over the underlying structure. We use these applications to assess the robustness of the FFT plug-in EM algorithm. These experiments provide a concrete picture of how the proposed approach behaves in finite samples and in practical settings. First we focus on the estimation of a single dominant component.
5.1 Simulated datasets
We first consider a mixture with a dominant component consisting of $60\%$ $N(10,1)$ and a contaminating component of $40\%$ $N(15,1)$. We generate $n=500$ observations from this model and run Algorithm 1 on the resulting sample. The estimated centre of the dominant population is $\hat{\mu}=9.98$, which is extremely close to the true centre $10$.
Figure 1: 60-40 mixture of N(10,1) and N(15,1)
Now we consider a mixture with a dominant component consisting of $60\%$ $N(10,16)$ – a diffuse and “spread-out" population and a relatively sharp contaminating component of $40\%$ $N(15,1)$. We sample $n=1000$ observations from this model and run Algorithm 1 on the resulting sample. The estimated centre of the dominant population is $\hat{\mu}=9.28$, which is remarkably close to the true centre $10$, despite the high variance of the primary population.
Figure 2: 60-40 mixture of N(10,16) and N(15,1)
Let us demonstrate our practical heuristic strategy now. Consider a mixture with a dominant component consisting of $60\%$ $N(10,1)$ and a diffuse contaminating component of $40\%$ $N(20,16)$. We again generate $n=1000$ observations from this model and run Algorithm 1 on the resulting sample. The nonparametric part in the vanilla weighted EM algorithm [37] absorbed the diffuse minority and estimated the centre of the dominant population as $\hat{\theta}=19.91$, which is actually close to the centre of the contaminating component. While our practical method ran two simultaneous EM algorithms and returned $\hat{\theta}=9.98$, which is indeed very close to the true location $10.$
Figure 3: Flipped estimates in a 60-40 mixture of N(10,1) and N(20,16)
Figure 4: Correct estimates in a 60-40 mixture of N(10,1) and N(20,16)
A runtime performance for this setup is summarized in Table 1. All simulations were run on a standard consumer-grade computer. Note that in all the runs, the vanilla EM was unable to identify the true dominant component.
Table 1: Runtime comparison (seconds) of FFT accelerated EM vs. vanilla EM.
Sample size FFT accelerated EM (s) Vanilla EM (s)
$n=500$ 0.049159 1.391762
$n=1000$ 0.043346 6.263643
$n=2000$ 0.059048 18.146281
$n=3000$ 0.059223 38.745702
$n=5000$ 0.113147 163.754303
$n=7000$ 0.132501 253.051952
$n=10000$ 0.342913 1145.869190
$n=12000$ 0.220001 2148.406933
$n=15000$ 0.315351 6035.796481
We now consider a mixture with a dominant component consisting of $60\%$ Cauchy with centre 0, scale $2$ and a contaminating component of $40\%$ Cauchy with centre $10$, scale $1$. We again generate $n=1000$ observations from this model. But now, we model the parametric part through a normal family in our working log-likelihood and run Algorithm 1 on the sample. The estimated centre of the dominant population is $\hat{\mu}=-0.25$, which is very close to the true centre 0, even after only approximately modeling the data through a Normal parametric family.
Figure 5: Cauchy mixture (60–40 at centers 0 & 10, scales 2 & 1)
Again, we consider a mixture with a dominant component consisting of $60\%$ Cauchy with centre 0, scale $5$ (diffuse majority) and a contaminating component of $40\%$ Cauchy with centre $10$, scale $1$ (tight minority). Generate $n=1000$ observations from this model. Once again, we model the parametric part through a normal family in the working log-likelihood and run Algorithm 1 on the sample. The estimated centre of the dominant population is $\hat{\mu}=1.08$, which is remarkably close to the true centre 0, even after only approximately modeling the data through a Normal family.
Figure 6: Cauchy mixture (60–40 at centers 0 & 10, scales 5 & 1)
5.2 Real datasets
This dataset records the lengths of Snapper fish, originally collected by [6]. The purpose of the data is to distinguish among the possible age groups to which an individual fish might belong—specifically, whether it comes from the current year’s spawning cohort or from that of the previous year. The empirical distribution is visibly bimodal, with modes located near 5 and 7.75 inches. The histogram shows a more pronounced peak around 5 inches, suggesting that fish from the current year’s spawning are represented in greater proportion than those from the earlier cohort.
Our algorithm estimates the average length of the dominant population to be $5.385$ inches, which is quite close to estimates calculated using more sophisticated techniques in other robust statistics literature.
Figure 7: Snapper data. Estimated mean 5.3856 inches
The second dataset is Newcomb’s classic light–speed measurements, previously analyzed by [2], [5], and others. The data records the time (in nanoseconds) required for light to traverse 7442 meters at sea level, subtracted from a baseline value of 24800 nanoseconds. The resulting histogram shows a prominent peak around 30, along with two well-known outliers at 2-2 and 44-44. We further shift the dataset by adding 44 to each observation so that all values become non-negative.
Clearly, the estimate from our algorithm is substantially close to the peak location of the larger population.
Figure 8: Newcomb light speed data. Estimated mean 71.922
This dataset contains the width–to–length ratios of $20$ beaded rectangles crafted by the Shoshoni Indians and used as decorative elements on clothing and blankets [21]. Our interest lies in estimating the central tendency of these ratios. In particular, one may ask whether the Shoshoni exhibited a preference for golden rectangles; see, for example, [22] for a discussion of golden rectangles. If such a preference existed, then the width–to–length ratios of their rectangles would be expected to cluster around the golden ratio, approximately $0.618$.
Our algorithm returns the centre of the dominant population as $0.63,$ quite close to $0.618$, suggesting empirical evidence in favour of the suspected preference of the Shoshoni people towards golden rectangles.
Figure 9: Ratio data. Estimated mean 0.630293
The fourth dataset consists of rainfall measurements (in millimeters) recorded every fourth day during the winter months (June, July, and August) of 1981–83 in Melbourne, Australia [30]. The dataset includes a substantial outlier of approximately 300 mm. For our analysis, we use a scale-transformed version of these observations.
The estimate from our algorithm is reasonably close to the peak location of the larger population, despite the presence of extreme outliers.
Figure 10: Rainfall data. Estimated mean 1.239 mm
Now let us assess the performance of our algorithm for sequential estimation of multiple significant peaks. As before, we run our algorithm on both simulated and real world datasets.
5.3 Simulated datasets
We first simulate $500$ samples from a $3$-component gaussian mixture $0.45\>\mathcal{N}(-6,1)+0.3\;\mathcal{N}(0,2)+0.25\;\mathcal{N}(8,2.5).$ Our sequential extraction scheme produces 3 peaks with centres 5.972,0.063, 7.9945.972,\;-0.063,\;7.994 which are very close to to the true peak locations.
Figure 11: 45-30-25 mixture of N(-6,1), N(0,2), N(8,2.5)
To examine a more difficult scenario, where the peaks are not very well separated and with considerable component overlap, we simulate $1000$ samples from a $3$-component gaussian mixture $0.3\>\mathcal{N}(-3,1.5)+0.4\;\mathcal{N}(0,1)+0.25\;\mathcal{N}(3,1.5).$ Our sequential extraction scheme produces 3 peaks with centres 2.866, 0.080, 2.651-2.866,\;0.080,\;2.651. Remarkably close to the true locations inspite of significant overlap between the individual components.
Figure 12: 30-40-30 mixture of N(-3,1.5), N(0,1), N(3,1.5)
Now let us see how our algorithm performs where there are even more components. We simulate $10000$ samples from a $5$-component gaussian mixture $0.15\;\mathcal{N}(-6,1)+0.20\;\mathcal{N}(-3,1)+0.25\;\mathcal{N}(0,1)+0.20\;\mathcal{N}(3,1)+0.20\;\mathcal{N}(6,1).$ Our algorithm produces 5 peaks with centres 6.184, 3.254, 0.013, 2.994, 6.089-6.184,\;3.254,\;0.013,\;2.994,\;6.089, which once again are sufficiently close to the true locations.
Figure 13: 15-20-25-20-20 mixture of N(-6,1), N(-3,1), N(0,1), N(3,1), N(6,1)
5.4 Real datasets
The classic Iris data set [12], contains 150 flowers with 50 observations from each of three species (Iris setosa, Iris versicolor, Iris virginica) and four measurements in centimetres (sepal length, sepal width, petal length, petal width); focusing on petal length, the species-wise empirical means and standard deviations are approximately $(\mu,\sigma)=(1.46,0.17)$ for setosa, $(4.26,0.47)$ for versicolor, and $(5.55,0.55)$ for virginica, so that the pooled marginal distribution of petal length is a trimodal mixture with three well-separated peaks corresponding to the three species. On applying our sequential algorithm on this dataset, we recover three distinct peaks with centers at $1.449$, $4.493$ and $5.600$, which are very good robust representatives of the empirical peak locations observed in nature, extracted from a relatively small sample.
Figure 14: Multiple component extraction in Iris dataset
We now return to the snapper dataset. In the current-year cohort there remains a substantial fraction of fish from previous spawning seasons, which are noticeably larger in size. Although the overall population has declined due to migration and fishing pressure, a non‑negligible proportion of these older fish is still present in the sample. Applying our sequential scheme to this dataset, we identified two pronounced modes at $5.342$ and $7.499$, corresponding respectively to the newer and older spawning groups.
Figure 15: Snapper data: 2 significant components
Proofs and supporting results
Proof of Lemma 1
Binning error.
Fix $y\in\mathbb{R}$ and set $g_{y}(t):=\mathbb{K}_{h_{n}}(y-t)$. Then $g_{y}\in C^{2}$ with $\|g_{y}^{\prime\prime}\|_{\infty}=\|\mathbb{K}^{\prime\prime}\|_{\infty}/h_{n}^{3}$. Since $Y_{i}=x_{j(i)}+\alpha_{i}\Delta_{n}\in[x_{j(i)},x_{j(i)+1}]$, the $C^{2}$ linear–interpolation remainder gives
$\big|g_{y}(Y_{i})-(1-\alpha_{i})g_{y}(x_{j(i)})-\alpha_{i}g_{y}(x_{j(i)+1})\big|\ \leq\ \frac{\Delta_{n}^{2}}{8}\,\frac{\|\mathbb{K}^{\prime\prime}\|_{\infty}}{h_{n}^{3}}.$
Multiplying by $\omega_{i}\geq 0$ and summing over $i$ (using $\sum_{i}\omega_{i}=1$) yields
$\big\|\tilde{f}^{\mathrm{bin}}_{1,n}-\tilde{f}^{\mathrm{exact}}_{1,n}\big\|_{\infty}\ \leq\ \frac{1}{8}\,\|\mathbb{K}^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
Exact FFT on grid.
With zero padding, the FFT computes the linear convolution on indices $0,\dots,M-1$, so $\tilde{f}^{\mathrm{FFT}}_{1,n}(x_{m})=\tilde{f}^{\mathrm{bin}}_{1,n}(x_{m})$ for all $m$.
Off–grid interpolation at $Y_{i}$.
Let $F(y):=\tilde{f}^{\mathrm{bin}}_{1,n}(y)=\sum_{m=0}^{M-1}c[m]\mathbb{K}_{h_{n}}(y-x_{m})$. Then $F\in C^{2}$ with
$F^{\prime\prime}(y)=\sum_{m=0}^{M-1}c[m]\,\mathbb{K}_{h_{n}}^{\prime\prime}(y-x_{m})\quad\Rightarrow\quad\|F^{\prime\prime}\|_{\infty}\leq\sum_{m}c[m]\|\mathbb{K}_{h_{n}}^{\prime\prime}\|_{\infty}=\frac{\|\mathbb{K}^{\prime\prime}\|_{\infty}}{h_{n}^{3}}.$
On $x_{j(i)},x_{j(i)+1}]$, the linear interpolant at $Y_{i}$ equals
$\widehat{F}(Y_{i})=(1-\alpha_{i})F(x_{j(i)})+\alpha_{i}F(x_{j(i)+1}),$
with interpolation error
$|F(Y_{i})-\widehat{F}(Y_{i})|\ \leq\ \frac{\Delta_{n}^{2}}{8}\,\frac{\|\mathbb{K}^{\prime\prime}\|_{\infty}}{h_{n}^{3}}.$
Because $\tilde{f}^{\mathrm{FFT}}_{1,n}(x_{m})=F(x_{m})$, our definition $\tilde{f}^{\mathrm{FFT}}_{1,n}(Y_{i})=\widehat{F}(Y_{i})$ gives
$|\tilde{f}^{\mathrm{FFT}}_{1,n}(Y_{i})-\tilde{f}^{\mathrm{bin}}_{1,n}(Y_{i})|\ \leq\ \frac{1}{8}\,\|\mathbb{K}^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
Triangle inequality.
Combine (B) and (E):
$|\tilde{f}^{\mathrm{FFT}}_{1,n}(Y_{i})-\tilde{f}^{\mathrm{exact}}_{1,n}(Y_{i})|\ \leq\ \frac{1}{8}\,\|\mathbb{K}^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}+\frac{1}{8}\,\|\mathbb{K}^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}\ =\ \frac{1}{4}\,\|\mathbb{K}^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
Taking $\max_{i}$ or $\sup_{y}$ gives the claim.
Proof of Proposition 1
Introduce latent indicators $Z_{i}\in\{0,1\}$ ($Z_{i}=1$ for the parametric component). The complete–data log–likelihood is
$\ell_{c}(\pi_{0},\theta,f_{1})=\sum_{i=1}^{n}\Big\{Z_{i}[\log\pi_{0}+\log f_{0}(Y_{i};\theta)]+(1-Z_{i})[\log(1-\pi_{0})+\log f_{1}(Y_{i})]\Big\}.$
Given $(\theta^{(t)},\pi_{0}^{(t)})$ and the current FFT background $\tilde{f}^{\mathrm{FFT},(t)}_{1,n}$, the E–step computes
$r_{i}^{(t)}=\mathbb{E}[Z_{i}\mid Y_{i}]=\frac{\pi_{0}^{(t)}f_{0}(Y_{i};\theta^{(t)})}{\pi_{0}^{(t)}f_{0}(Y_{i};\theta^{(t)})+(1-\pi_{0}^{(t)})\tilde{f}^{\mathrm{FFT},(t)}_{1,n}(Y_{i})},\quad s_{i}^{(t)}=1-r_{i}^{(t)}.$
Taking conditional expectation of $\ell_{c}$ yields
$Q^{(t)}(\pi_{0},\theta,f_{1})=\frac{1}{n}\sum_{i=1}^{n}\Big\{r_{i}^{(t)}\,[\log\pi_{0}+\log f_{0}(Y_{i};\theta)]+s_{i}^{(t)}\,[\log(1-\pi_{0})+\log f_{1}(Y_{i})]\Big\}.$
Maximizing $Q^{(t)}$ in $\pi_{0}\in(0,1)$ with $\pi_{1}=1-\pi_{0}$ gives $\sum_{i}\big(\frac{r_{i}^{(t)}}{\pi_{0}}-\frac{s_{i}^{(t)}}{1-\pi_{0}}\big)=0$, hence $\pi_{0}^{(t+1)}=\frac{1}{n}\sum_{i}r_{i}^{(t)}$. Maximizing in $\theta$ gives the weighted MLE $\theta^{(t+1)}\in\arg\max_{\theta}\frac{1}{n}\sum_{i}r_{i}^{(t)}\log f_{0}(Y_{i};\theta)$, equivalently the weighted score equation stated.
For the nonparametric part, we adopt the sieve class $\mathcal{F}_{h_{n},\Delta_{n}}=\{y\mapsto\sum_{m=0}^{M-1}c[m]K_{h_{n}}(y-x_{m}):c[m]\geq 0,\sum_{m}c[m]=1\}$ and construct the next background by smoothing the normalized nonparametric responsibilities: first map $\{\omega_{i}^{(t)}\}$ to grid coefficients $c^{(t)}[\cdot]$ via linear binning, then compute the grid convolution by zero–padded FFT, and finally use linear interpolation at $Y_{i}$ in the next E–step. This completes the EM iteration with the FFT plug–in background.
Consistency proof and supporting lemmas
Assumption 1.
  1. (i)
    $f(y):=\pi_{0}f_{0}(y;\theta)+\pi_{1}f_{1}(y)$ satisfies $f(y)>0$ for all $y\in\mathbb{R}$.
  2. (ii)
    $(y,\theta)\mapsto f_{0}(y;\theta)$ is continuous on $\mathbb{R}\times\mathcal{C}$.
  3. (iii)
    For every $M<\infty$ there exists $c(M)>0$ such that
    $\inf_{\begin{subarray}{c}|y|\leq M\\ \theta\in\mathcal{C}\end{subarray}}f_{0}(y;\theta)\;\geq\;c(M).$
  4. (iv)
    There exists $B<\infty$ with $\sup_{\theta\in\mathcal{C},\,y}f_{0}(y;\theta)\leq B$ and $\sup_{y}f_{1}^{*}(y)\leq B.$
  5. (v)
    The population log–likelihood
    $L(\theta):=\mathbb{E}\Big[\log\{\pi_{0}f_{0}(Y;\theta)+\pi_{1}f_{1}(Y)\}\Big],\qquad\theta\in\mathcal{C},$
    is continuous on $\mathcal{C}$ and uniquely maximised at $\theta^{*}$.
  6. (vi)
    There exists $H:\mathbb{R}\to[0,\infty)$ with $\mathbb{E}[H(Y)]<\infty$ such that
    $\sup_{\theta\in\mathcal{C}}\big|\log\{\pi_{0}f_{0}(Y;\theta)+\pi_{1}f_{1}^{*}(Y)\}\big|\;\leq\;H(Y)\quad a.s.$
Assumption 2.
$\mathbb{K}$ is even, bounded, Lipschitz and $C^{2}$, with
$\int\mathbb{K}=1,\qquad\int u\mathbb{K}(u)\,du=0,\qquad\int u^{2}|\mathbb{K}(u)|\,du<\infty,\qquad\|\mathbb{K}^{\prime\prime}\|_{\infty}<\infty.$
Let $K_{h}(u):=h^{-1}K(u/h)$ and assume
$h_{n}\downarrow 0,\qquad\frac{nh_{n}}{\log(1/h_{n})}\ \longrightarrow\ \infty.$
For the FFT/binned implementation, use a uniform grid of spacing $\Delta_{n}\downarrow 0$ and define
$\varepsilon_{n}\;:=\;\frac{\|\mathbb{K}^{\prime\prime}\|_{\infty}}{4}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}\ \longrightarrow\ 0.$
Let $\tilde{f}_{1,n}^{\mathrm{exact}}$ be the weighted KDE and $\tilde{f}_{1,n}^{\mathrm{FFT}}$ its binned/FFT approximation. Assume the standard $C^{2}$ bound
$\sup_{y\in\mathbb{R}}\big|\tilde{f}_{1,n}^{\mathrm{FFT}}(y)-\tilde{f}_{1,n}^{\mathrm{exact}}(y)\big|\;\leq\;\varepsilon_{n}.$
Assumption 3.
There exist nonnegative (possibly data–dependent) weights $\hat{\omega}_{1},\dots,\hat{\omega}_{n}$ with $\sum_{i=1}^{n}\hat{\omega}_{i}=1$ such that
$\tilde{f}_{1,n}^{\mathrm{exact}}(y)\;:=\;\sum_{i=1}^{n}\hat{\omega}_{i}\,\mathbb{K}_{h_{n}}(y-Y_{i}),\qquad y\in\mathbb{R}.$
There exists a deterministic $D_{n}\uparrow\infty$ for which
$\sup_{|y|\leq D_{n}}\big|\tilde{f}_{1,n}^{\mathrm{exact}}(y)-f_{1}^{*}(y)\big|\;=\;O_{p}\!\Big(h_{n}^{2}+\sqrt{\tfrac{\log(1/h_{n})}{nh_{n}}}\Big),$
and hence, by Assumption 2,
$\sup_{|y|\leq D_{n}}\big|\tilde{f}_{1,n}^{\mathrm{FFT}}(y)-f_{1}^{*}(y)\big|\;=\;O_{p}\!\Big(h_{n}^{2}+\sqrt{\tfrac{\log(1/h_{n})}{nh_{n}}}+\varepsilon_{n}\Big).$
Remark. Such an assumption is common in the literature, for example – [17], [14], [11], [33], [20].
For convenience, write
$r_{n}\;:=\;h_{n}^{2}+\sqrt{\frac{\log(1/h_{n})}{nh_{n}}}+\varepsilon_{n}.$
Assumption 4.
For the $D_{n}$ in Assumption 3,
$\frac{r_{n}}{c(D_{n})}\ \longrightarrow\ 0,\qquad\Pr(|Y|>D_{n})\,\log(1/h_{n})\ \longrightarrow\ 0.$
Assumption 5.
There exists an $H_{0}:\mathbb{R}\to[0,\infty)$ with $\mathbb{E}[H_{0}(Y)]<\infty$ such that
$\sup_{\theta\in\mathcal{C}}\big|\log f_{0}(Y;\theta)\big|\;\leq\;H_{0}(Y)\quad a.s.$
Lemma 2.
Under Assumption 1,
$\sup_{\theta\in\mathcal{C}}\left|\frac{1}{n}\sum_{i=1}^{n}\log\{\pi_{0}f_{0}(Y_{i};\theta)+\pi_{1}f_{1}^{*}(Y_{i})\}-\mathcal{L}(\theta)\right|\ \xrightarrow{p}\ 0.$
Proof.
For each $y$, $\theta\mapsto\log\{\pi_{0}f_{0}(y;\theta)+\pi_{1}f_{1}^{*}(y)\}$ is continuous on compact $\mathcal{C}$ and dominated by $H(Y)$ with $\mathbb{E}H(Y)<\infty$; see [34, Thm 19.4]. ∎
Proposition 2 (Uniform convergence of the FFT plug–in criterion).
We work under Assumptions 1, 2, 3, 4 and 5. Let
$\tilde{\mathcal{L}_{n}}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\log\!\Big\{\pi_{0}f_{0}(Y_{i};\theta)+\pi_{1}\,\tilde{f}_{1,n}^{\mathrm{FFT}}(Y_{i})\Big\},\qquad\theta\in\mathcal{C},$
and $L(\theta):=\mathbb{E}\log\{\pi_{0}f_{0}(Y;\theta)+\pi_{1}f_{1}^{*}(Y)\}.$ Then
$\sup_{\theta\in\mathcal{C}}\big|\tilde{\mathcal{L}}_{n}(\theta)-\mathcal{L}(\theta)\big|\ \xrightarrow{p}\ 0.$
Proof.
Write $\ell_{\theta}^{u}(y):=\log\{\pi_{0}f_{0}(y;\theta)+\pi_{1}u(y)\}$, and decompose
$\tilde{L}_{n}(\theta)-L(\theta)=R_{1,n}(\theta)+R_{2,n}(\theta),\qquad R_{1,n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\big(\ell_{\theta}^{\tilde{f}}(Y_{i})-\ell_{\theta}^{f_{1}^{*}}(Y_{i})\big),\quad R_{2,n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{f_{1}^{*}}(Y_{i})-L(\theta),$
with $\tilde{f}:=\tilde{f}_{1,n}^{\mathrm{FFT}}$. By Lemma 2 (oracle ULLN), $\sup_{\theta}|R_{2,n}(\theta)|\to_{p}0$. We control $R_{1,n}$ via a body/tail split at the deterministic window $M_{n}\uparrow\infty$ from Assumption 4:
$R_{1,n}(\theta)=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\big(\ell_{\theta}^{\tilde{f}}(Y_{i})-\ell_{\theta}^{f_{1}^{*}}(Y_{i})\big)\mathbb{I}_{\{|Y_{i}|\leq M_{n}\}}}_{=:A_{n,M_{n}}(\theta)}+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\big(\ell_{\theta}^{\tilde{f}}(Y_{i})-\ell_{\theta}^{f_{1}^{*}}(Y_{i})\big)\mathbb{I}_{\{|Y_{i}|>M_{n}\}}}_{=:B_{n,M_{n}}(\theta)}.$
Body: On $\{|Y_{i}|\leq M_{n}\}$, Assumption 1(iii) gives $\pi_{0}f_{0}(Y_{i};\theta)\geq\pi_{0}c(M_{n})$, hence by the mean value theorem for $\log$,
$\big|\ell_{\theta}^{\tilde{f}}(Y_{i})-\ell_{\theta}^{f_{1}^{*}}(Y_{i})\big|\leq\frac{\pi_{1}}{\pi_{0}c(M_{n})}\,\big|\tilde{f}(Y_{i})-f_{1}^{*}(Y_{i})\big|.$
Therefore, using the trivial averaging inequality $\frac{1}{n}\sum_{i=1}^{n}|a_{i}|\mathbb{I}_{\{|Y_{i}|\leq M_{n}\}}\leq\sup_{|y|\leq M_{n}}|a(y)|$ with $a(y)=\tilde{f}(y)-f_{1}^{*}(y)$, we obtain
$\sup_{\theta\in\mathcal{C}}|A_{n,M_{n}}(\theta)|\leq\frac{\pi_{1}}{\pi_{0}c(M_{n})}\,\sup_{|y|\leq M_{n}}\big|\tilde{f}(y)-f_{1}^{*}(y)\big|=O_{p}\!\Big(\frac{r_{n}}{c(M_{n})}\Big)=o_{p}(1),$
by Assumptions 3 and 4.
Tail: Fix $\theta$ and $y$. Since $\tilde{f}(y)\geq 0$, we have the upper bound
$\ell_{\theta}^{\tilde{f}}(y)=\log\{\pi_{0}f_{0}(y;\theta)+\pi_{1}\tilde{f}(y)\}\leq\log\{\pi_{0}B+\pi_{1}(\|\mathbb{K}\|_{\infty}h_{n}^{-1}+\varepsilon_{n})\}\leq C_{0}+\log(1/h_{n})$
for all large $n$ by Assumptions 1(iv) and 2. We also have the lower bound $\ell_{\theta}^{\tilde{f}}(y)\geq\log\{\pi_{0}f_{0}(y;\theta)\}.$ Combining,
$\big|\ell_{\theta}^{\tilde{f}}(y)\big|\leq\max\!\big\{C_{0}+\log(1/h_{n}),\,-\log(\pi_{0}f_{0}(y;\theta))\big\}\leq C_{1}+\log(1/h_{n})+\big|\log f_{0}(y;\theta)\big|,$
with $C_{1}$ absorbing $C_{0}$ and $|\log\pi_{0}|$. By Assumption 5, $\sup_{\theta\in\mathcal{C}}|\log f_{0}(Y;\theta)|\leq H_{0}(Y)$ and hence
$\sup_{\theta\in\mathcal{C}}\big|\ell_{\theta}^{\tilde{f}}(Y)\big|\leq C_{1}+\log(1/h_{n})+H_{0}(Y).$
For the oracle term, Assumption 1 gives $\sup_{\theta\in\mathcal{C}}\big|\ell_{\theta}^{f_{1}^{*}}(Y)\big|\leq H(Y).$ Therefore,
$\sup_{\theta\in\mathcal{C}}|B_{n,M_{n}}(\theta)|\leq\frac{1}{n}\sum_{i=1}^{n}\Big(C_{1}+\log(1/h_{n})+H_{0}(Y_{i})+H(Y_{i})\Big)\mathbb{I}_{\{|Y_{i}|>M_{n}\}}.$
Take expectations and use LLN and Markov inequality:
$\mathbb{E}\!\left[\frac{1}{n}\sum_{i=1}^{n}(H_{0}(Y_{i})+H(Y_{i}))\mathbb{I}_{\{|Y_{i}|>M_{n}\}}\right]=\mathbb{E}[(H_{0}(Y)+H(Y))\mathbb{I}_{\{|Y|>M_{n}\}}]\;\longrightarrow\;0$
by dominated convergence since $H_{0},H\in L^{1}(g)$ and $M_{n}\uparrow\infty$. Also, by Assumption 4, $\Pr(|Y|>M_{n})\log(1/h_{n})\to 0.$ As $n^{-1}\sum_{i=1}^{n}\mathbb{I}_{\{|Y_{i}|>M_{n}\}}=P_{n}(|Y|>M_{n})=\Pr(|Y|>M_{n})+o_{p}(1)$, we conclude
$\sup_{\theta\in\mathcal{C}}|B_{n,M_{n}}(\theta)|=o_{p}(1).$
Combining the body and tail bounds, $\sup_{\theta\in\mathcal{C}}|R_{1,n}(\theta)|\to_{p}0.$ Together with the oracle ULLN for $R_{2,n}$ this yields $\sup_{\theta\in\mathcal{C}}|\tilde{\mathcal{L}_{n}}(\theta)-\mathcal{L}(\theta)|\to_{p}0.$
Proof of Theorem 2 (Consistency)
By Proposition 2, $\sup_{\theta\in\mathcal{C}}|\tilde{\mathcal{L}}_{n}(\theta)-\mathcal{L}(\theta)|\to_{p}0.$ Assumption 1(v) yields uniqueness and continuity of the maximiser on compact $\mathcal{C}$. Now apply Wald’s argmax theorem.
CLT assumptions and supporting lemmas
Write
$\ell(y;\theta):=\log\{\pi_{0}f_{0}(y;\theta)+\pi_{1}f_{1}^{*}(y)\},\qquad\dot{\ell}(y;\theta):=\nabla_{\theta}\ell(y;\theta),\qquad\ddot{\ell}(y;\theta):=\nabla_{\theta}^{2}\ell(y;\theta),$
and
$S_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\dot{\ell}(Y_{i};\theta),\qquad H_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\ddot{\ell}(Y_{i};\theta).$
We have
$\tilde{\mathcal{L}}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\log\!\Big\{\pi_{0}f_{0}(Y_{i};\theta)+\pi_{1}\,\tilde{f}_{1,n}^{\mathrm{FFT}}(Y_{i})\Big\},\qquad\theta\in\mathcal{C},$
with maximiser $\hat{\theta}_{n}\in\arg\max_{\theta\in\mathcal{C}}\tilde{L}_{n}(\theta)$. By Theorem 2, $\hat{\theta}_{n}\to_{p}\theta^{*}$.
Recall
$r_{n}:=h_{n}^{2}+\sqrt{\frac{\log(1/h_{n})}{nh_{n}}}+\varepsilon_{n},$
and the window $D_{n}\uparrow\infty$ from Assumption 4.
Assumption 6.
For each $y$, $\theta\mapsto f_{0}(y;\theta)$ is twice continuously differentiable on $\mathcal{C}$, and differentiation under the integral sign is valid for $\ell,\dot{\ell},\ddot{\ell}$. There exist envelopes $H_{1},H_{2}\in L^{2}(f)$ such that, for all $\theta\in\mathcal{C}$,
$\|\dot{\ell}(Y;\theta)\|\leq H_{1}(Y),\qquad\|\ddot{\ell}(Y;\theta)\|\leq H_{2}(Y).$
Moreover, for each $M<\infty$,
$\sup_{\begin{subarray}{c}|y|\leq M\\ \theta\in\mathcal{C}\end{subarray}}\Big(\|\nabla_{\theta}f_{0}(y;\theta)\|+\|\nabla_{\theta}^{2}f_{0}(y;\theta)\|\Big)<\infty.$
The Fisher information
$I(\theta^{*}):=\mathbb{E}_{f}\!\big[\dot{\ell}(Y;\theta^{*})\dot{\ell}(Y;\theta^{*})^{\top}\big]=-\,\mathbb{E}_{f}\!\big[\ddot{\ell}(Y;\theta^{*})\big]$
exists, is finite, and is positive definite.
Assumption 7.
There exist measurable functions $H_{1,n},H_{2,n}:\mathbb{R}\to[0,\infty)$ such that, for all $\theta\in\mathcal{C}$,
$\Big\|\nabla_{\theta}\log\{\pi_{0}f_{0}(Y;\theta)+\pi_{1}\tilde{f}_{1,n}^{\mathrm{FFT}}(Y)\}\Big\|\leq H_{1,n}(Y),$
$\Big\|\nabla_{\theta}^{2}\log\{\pi_{0}f_{0}(Y;\theta)+\pi_{1}\tilde{f}_{1,n}^{\mathrm{FFT}}(Y)\}\Big\|\leq H_{2,n}(Y),$
and, with $D_{n}$ as in Assumption 4,
$\sqrt{n}\,\mathbb{E}_{f}\!\big[(H_{1,n}(Y)+H_{1}(Y))\,\mathbb{I}_{\{|Y|>D_{n}\}}\big]\ \longrightarrow\ 0,\qquad\mathbb{E}_{f}\!\big[(H_{2,n}(Y)+H_{2}(Y))\,\mathbb{I}_{\{|Y|>D_{n}\}}\big]\ \longrightarrow\ 0.$
Assumption 8.
Let
$\Phi(y):=\frac{\pi_{0}\pi_{1}\,\nabla_{\theta}f_{0}(y;\theta^{*})}{\{\pi_{0}f_{0}(y;\theta^{*})+\pi_{1}f_{1}^{*}(y)\}^{2}}\in\mathbb{R},$
and define the windowed linear functional
$\Lambda_{n}:=\frac{1}{n}\sum_{i=1}^{n}\Phi(Y_{i})\,\big(\tilde{f}_{1,n}^{\mathrm{FFT}}(Y_{i})-f_{1}^{*}(Y_{i})\big)\,\mathbb{I}_{\{|Y_{i}|\leq D_{n}\}}\in\mathbb{R}.$
Then $\sqrt{n}\,\|\Lambda_{n}\|\to_{p}0$.
Assumption 9.
With $r_{n}$ and $D_{n}$ as above,
$\frac{\sqrt{n}\,r_{n}^{2}}{c(D_{n})^{3}}\ \longrightarrow\ 0,$
where $c(M)$ is from Assumption 1(iii).
Lemma 3 (Oracle score CLT).
Under Assumptions 1, and 6,
$\sqrt{n}\,S_{n}(\theta^{*})\ \rightsquigarrow\ \mathcal{N}\big(0,\,I(\theta^{*})\big).$
Proof.
$\dot{\ell}(Y_{i};\theta^{*})$ are i.i.d. with mean 0, covariance $I(\theta^{*})$, and $\mathbb{E}\|\dot{\ell}(Y;\theta^{*})\|^{2}\leq\mathbb{E}[H_{1}(Y)^{2}]<\infty$. Apply the multivariate Lindeberg–Feller CLT. ∎
Lemma 4.
Under Assumptions 1 and  6,
$\sup_{\theta\in\mathcal{C}}\big\|H_{n}(\theta)-\mathbb{E}_{f}[\ddot{\ell}(Y;\theta)]\big\|\ \xrightarrow{p}\ 0.$
In particular, by continuity of $\theta\mapsto\mathbb{E}_{f}[\ddot{\ell}(Y;\theta)]$ and Assumption 6, we have $H_{n}(\theta^{*})\to_{p}-I(\theta^{*})$.
Proof.
For each $y$, $\theta\mapsto\ddot{\ell}(y;\theta)$ is continuous on compact $\mathcal{C}$, and $\|\ddot{\ell}(Y;\theta)\|\leq H_{2}(Y)$ with $\mathbb{E}H_{2}(Y)<\infty$. A dominated Glivenko–Cantelli theorem for parametric classes (e.g. [34, Thm. 19.4]) yields the uniform LLN. Continuity of $\theta\mapsto\mathbb{E}[\ddot{\ell}(Y;\theta)]$ then gives the specialization at $\theta^{*}$. ∎
Lemma 5.
Let
$\tilde{S}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}\log\{\pi_{0}f_{0}(Y_{i};\theta)+\pi_{1}\tilde{f}_{1,n}^{\mathrm{FFT}}(Y_{i})\},\quad\tilde{H}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}^{2}\log\{\pi_{0}f_{0}(Y_{i};\theta)+\pi_{1}\tilde{f}_{1,n}^{\mathrm{FFT}}(Y_{i})\}.$
Under Assumptions 1, 2, 3, 4, 6, 7, 8 and 9,
$\sqrt{n}\,\big\|\tilde{S}_{n}(\theta^{*})-S_{n}(\theta^{*})\big\|\ \xrightarrow{p}\ 0,\qquad\sup_{\theta\in\mathcal{C}}\big\|\tilde{H}_{n}(\theta)-H_{n}(\theta)\big\|\ \xrightarrow{p}\ 0.$
Proof.
Write $\tilde{s}_{i}(\theta)$ and $s_{i}(\theta)$ for the summand scores, and split with the window $D_{n}$:
$\tilde{S}_{n}(\theta^{*})-S_{n}(\theta^{*})=\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\tilde{s}_{i}-s_{i})\,\mathbb{I}_{\{|Y_{i}|\leq D_{n}\}}}_{=:B_{n}}+\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\tilde{s}_{i}-s_{i})\,\mathbb{I}_{\{|Y_{i}|>D_{n}\}}}_{=:T_{n}}.$
By Assumption 7, $\|\tilde{s}_{i}-s_{i}\|\leq H_{1,n}(Y_{i})+H_{1}(Y_{i})$, so
$\sqrt{n}\,\|T_{n}\|\leq\sqrt{n}\,\frac{1}{n}\sum_{i=1}^{n}(H_{1,n}(Y_{i})+H_{1}(Y_{i}))\,\mathbb{I}_{\{|Y_{i}|>D_{n}\}}.$
Define
$Z_{n}:=\sqrt{n}\,\frac{1}{n}\sum_{i=1}^{n}(H_{1,n}(Y_{i})+H_{1}(Y_{i}))\,\mathbb{I}_{\{|Y_{i}|>D_{n}\}}.$
Then $\mathbb{E}[Z_{n}]=\sqrt{n}\,\mathbb{E}[(H_{1,n}+H_{1})\mathbb{I}_{\{|Y|>D_{n}\}}]\to 0$ by Assumption 7, so $Z_{n}=o_{p}(1)$ by Markov. Thus $\sqrt{n}\,\|T_{n}\|=o_{p}(1)$. The Hessian tail is analogous with $H_{2,n},H_{2}$ and no $\sqrt{n}$ factor.
For $|Y_{i}|\leq D_{n}$, the map $u\mapsto\nabla_{\theta}\log\{\pi_{0}f_{0}(Y_{i};\theta^{*})+\pi_{1}u(Y_{i})\}$ is twice Fréchet–differentiable near $u=f_{1}^{*}$ with first derivative $D_{u}[\nabla_{\theta}\log](y)\cdot v=-\Phi(y)\,v(y)$ and second derivative bounded, on $\{|y|\leq D_{n}\}$, by $C/c(D_{n})^{3}$. With $\Delta_{n}(y):=\tilde{f}_{1,n}^{\mathrm{FFT}}(y)-f_{1}^{*}(y)$,
$\tilde{s}_{i}(\theta^{*})-s_{i}(\theta^{*})=-\Phi(Y_{i})\,\Delta_{n}(Y_{i})+R_{i},\qquad\|R_{i}\|\lesssim\frac{\Delta_{n}(Y_{i})^{2}}{c(D_{n})^{3}}.$
Thus
$B_{n}=-\,\Lambda_{n}+\frac{1}{n}\sum_{i=1}^{n}R_{i}\,\mathbb{I}_{\{|Y_{i}|\leq D_{n}\}}.$
By Assumption 8, $\sqrt{n}\,\|\Lambda_{n}\|\to_{p}0$. By Assumption 3, $\sup_{|y|\leq D_{n}}|\Delta_{n}(y)|=O_{p}(r_{n})$, hence
$\sqrt{n}\cdot\frac{1}{n}\sum_{i=1}^{n}\|R_{i}\|\,\mathbb{I}_{\{|Y_{i}|\leq D_{n}\}}=O_{p}\!\Big(\frac{\sqrt{n}\,r_{n}^{2}}{c(D_{n})^{3}}\Big)=o_{p}(1)$
by Assumption 9. This proves the score claim. The Hessian claim is obtained by an analogous expansion with second derivatives. ∎
Proof of Theorem 3 (CLT)
Consistency gives $\hat{\theta}_{n}\to_{p}\theta^{*}$. The first–order condition $\tilde{S}_{n}(\hat{\theta}_{n})=0$ and a Taylor expansion imply
$0=\tilde{S}_{n}(\theta^{*})+\tilde{H}_{n}(\tilde{\theta}_{n})(\hat{\theta}_{n}-\theta^{*}),$
so
$\sqrt{n}(\hat{\theta}_{n}-\theta^{*})=-\,\tilde{H}_{n}(\tilde{\theta}_{n})^{-1}\,\sqrt{n}\,\tilde{S}_{n}(\theta^{*}).$
Decompose
$\sqrt{n}\,\tilde{S}_{n}(\theta^{*})=\sqrt{n}\,S_{n}(\theta^{*})+\sqrt{n}\,(\tilde{S}_{n}(\theta^{*})-S_{n}(\theta^{*})).$
By Lemma 3,
$\sqrt{n}\,S_{n}(\theta^{*})\Rightarrow\mathcal{N}(0,I(\theta^{*}))$
By Lemma 5,
$\sqrt{n}\,(\tilde{S}_{n}(\theta^{*})-S_{n}(\theta^{*}))=o_{p}(1)$
and
$\sup_{\theta}\|\tilde{H}_{n}(\theta)-H_{n}(\theta)\|\to_{p}0$
By Lemma 4 and the continuity of $\theta\mapsto\mathbb{E}[\ddot{\ell}(Y;\theta)]$, we also have $H_{n}(\theta)\to_{p}\mathbb{E}[\ddot{\ell}(Y;\theta)]$ uniformly and, with $\tilde{\theta}_{n}\to_{p}\theta^{*}$, we have $\tilde{H}_{n}(\tilde{\theta}_{n})\to_{p}-I(\theta^{*}).$ Now Slutsky’s theorem yields the claim.
Consistency proof and supporting lemmas (sequential scheme)
Assumption 10.
  1. (i)
    $f_{0}(\cdot;\theta)$ is a Lebesgue density for each $\theta\in\mathcal{C}$, $f_{\mathrm{bg}}^{(\ell)}$ is a Lebesgue density, and $f^{(\ell)}(y)>0$ for all $y$.
  2. (ii)
    $(y,\theta)\mapsto f_{0}(y;\theta)$ is continuous on $\mathbb{R}\times\mathcal{C}$.
  3. (iii)
    For every $M<\infty$ there exists $c_{\ell}(M)>0$ such that
    $\inf_{\begin{subarray}{c}|y|\leq M\\ \theta\in\mathcal{C}\end{subarray}}f_{0}(y;\theta)\geq c_{\ell}(M).$
  4. (iv)
    There exists $B_{\ell}<\infty$ with $\sup_{\theta\in\mathcal{C},\,y}f_{0}(y;\theta)\leq B_{\ell}$ and $\sup_{y}f_{\mathrm{bg}}^{(\ell)}(y)\leq B_{\ell}$.
  5. (v)
    The oracle criterion $L^{(\ell)}$ is continuous on $\mathcal{C}\times\Pi_{\min}$ and uniquely maximised at $\eta_{\ell}=(\theta_{\ell},\alpha_{\ell}^{(\ell)})$.
  6. (vi)
    (Oracle log–envelope.) There exists $H^{(\ell)}\in L^{1}(f^{(\ell)})$ such that
    $\sup_{\theta\in\mathcal{C},\;\alpha\in\Pi_{\min}}\big|\ell^{(\ell)}\big(Y;\eta,f_{\mathrm{bg}}^{(\ell)}\big)\big|\leq H^{(\ell)}(Y)\quad a.s.$
Assumption 11.
Let $K$ be even, bounded, Lipschitz and $C^{2}$ with $\int K=1$, $\int uK(u)\,du=0$, $\int u^{2}|K(u)|\,du<\infty$ and $\|K^{\prime\prime}\|_{\infty}<\infty$. Define $K_{h}(u):=h_{n}^{-1}K(u/h_{n})$ and assume $h_{n}\downarrow 0$, $nh_{n}/\log(1/h_{n})\to\infty$. Let $\tilde{f}_{\ell,n}^{\mathrm{exact}}$ denote the exact weighted KDE with weights $\widehat{w}_{i}^{(\ell)}$:
$\tilde{f}_{\ell,n}^{\mathrm{exact}}(y):=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}K_{h_{n}}(y-Y_{i}),$
and $\tilde{f}_{\ell,n}^{\mathrm{FFT}}$ its binned/FFT approximation on a grid of spacing $\Delta_{n}\downarrow 0$. Suppose there exists a deterministic window $M_{n}\uparrow\infty$ such that
$\sup_{|y|\leq M_{n}}\big|\tilde{f}_{\ell,n}^{\mathrm{exact}}(y)-f_{\mathrm{bg}}^{(\ell),*}(y)\big|=O_{p}\!\Big(h_{n}^{2}+\sqrt{\tfrac{\log(1/h_{n})}{nh_{n}}}\Big),$
and
$\sup_{y\in\mathbb{R}}\big|\tilde{f}_{\ell,n}^{\mathrm{FFT}}(y)-\tilde{f}_{\ell,n}^{\mathrm{exact}}(y)\big|\leq\varepsilon_{n},\qquad\varepsilon_{n}\to 0.$
In particular, for
$r_{n}:=h_{n}^{2}+\sqrt{\tfrac{\log(1/h_{n})}{nh_{n}}}+\varepsilon_{n},$
we have
$\sup_{|y|\leq M_{n}}\big|\tilde{f}_{\ell,n}^{\mathrm{FFT}}(y)-f_{\mathrm{bg}}^{(\ell),*}(y)\big|=O_{p}(r_{n}).$
Assumption 12.
With $M_{n}$ and $r_{n}$ as in Assumption 11, we have
$\frac{r_{n}}{c_{\ell}(M_{n})}\ \longrightarrow\ 0,\qquad\Pr_{f^{(\ell)}}(|Y|>M_{n})\,\log(1/h_{n})\ \longrightarrow\ 0.$
Assumption 13.
  1. (i)
    For every measurable $\Phi:\mathbb{R}\to\mathbb{R}$ with $\mathbb{E}_{f^{(\ell)}}|\Phi(Y)|<\infty$,
    $P_{n}^{(\ell)}\Phi:=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}\Phi(Y_{i})\ \xrightarrow{p}\ \mathbb{E}_{f^{(\ell)}}[\Phi(Y)].$
  2. (ii)
    There exists $C_{\ell}<\infty$ such that $0\leq\widehat{w}_{i}^{(\ell)}\leq C_{\ell}/n$ for all $i$ and all large $n$.
Assumption 14.
There exists $H^{(\ell)}_{\mathrm{plug}}\in L^{1}(f^{(\ell)})$ such that, for all $\eta\in\mathcal{C}\times\Pi_{\min}$,
$\big|\ell^{(\ell)}\big(Y;\eta,\tilde{f}_{\ell,n}^{\mathrm{FFT}}\big)\big|\leq H^{(\ell)}_{\mathrm{plug}}(Y)\quad a.s.\ \text{for all large $n$},$
and $\mathbb{E}_{f^{(\ell)}}[H^{(\ell)}_{\mathrm{plug}}(Y)\,\mathbf{1}_{\{|Y|>M_{n}\}}]\to 0.$
[…]
Under Assumptions 1014,
$\sup_{\eta\in\mathcal{C}\times\Pi_{\min}}\big|\tilde{\mathcal{L}}_{n}^{(\ell)}(\eta)-\mathcal{L}^{(\ell)}(\eta)\big|\ \xrightarrow{p}\ 0.$
Proof.
Fix $\eta=(\theta,\alpha)$ and abbreviate $\ell_{\eta}^{u}(y):=\ell^{(\ell)}(y;\eta,u).$ Write
$\tilde{\mathcal{L}}_{n}^{(\ell)}(\eta)-\mathcal{L}^{(\ell)}(\eta)=R_{1,n}(\eta)+R_{2,n}(\eta),$
where
$R_{1,n}(\eta):=P_{n}^{(\ell)}\big(\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}-\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}\big),\quad R_{2,n}(\eta):=P_{n}^{(\ell)}\big(\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}\big)-\mathbb{E}_{f^{(\ell)}}\big[\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(Y)\big].$
Control of $R_{1,n}$. Let $M_{n}$ be as in Assumption 11 and write $R_{1,n}=A_{n,M_{n}}+B_{n,M_{n}}$ with
$A_{n,M_{n}}(\eta):=P_{n}^{(\ell)}\big(\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}-\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}\big)\mathbf{1}_{\{|Y|\leq M_{n}\}},$
$B_{n,M_{n}}(\eta):=P_{n}^{(\ell)}\big(\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}-\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}\big)\mathbf{1}_{\{|Y|>M_{n}\}}.$
On $\{|Y|\leq M_{n}\}$, Assumption 10(iii) and $\alpha\in\Pi_{\min}$ imply $\alpha f_{0}(Y;\theta)\geq\underline{\alpha}_{\ell},c_{\ell}(M_{n}).$ By the mean value theorem,
$\big|\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}(Y)-\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(Y)\big|\leq\frac{1-\underline{\alpha}_{\ell}}{\underline{\alpha}_{\ell},c_{\ell}(M_{n})}\big|\tilde{f}_{\ell,n}^{\mathrm{FFT}}(Y)-f_{\mathrm{bg}}^{(\ell),*}(Y)\big|$
on $\{|Y|\leq M_{n}\}$. Using the weighted analogue of the trivial averaging inequality,
$\Big|\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}\Delta(Y_{i})\mathbf{1}_{\{|Y_{i}|\leq M_{n}\}}\Big|\leq\sup_{|y|\leq M_{n}}|\Delta(y)|,$
with $\Delta(y)=\tilde{f}_{\ell,n}^{\mathrm{FFT}}(y)-f_{\mathrm{bg}}^{(\ell),*}(y)$, we obtain
$\sup_{\eta}|A_{n,M_{n}}(\eta)|\leq\frac{1-\underline{\alpha}_{\ell}}{\underline{\alpha}_{\ell},c_{\ell}(M_{n})}\sup_{|y|\leq M_{n}}\big|\tilde{f}_{\ell,n}^{\mathrm{FFT}}(y)-f_{\mathrm{bg}}^{(\ell),*}(y)\big|=O_{p}\!\Big(\frac{r_{n}}{c_{\ell}(M_{n})}\Big)=o_{p}(1),$
by Assumptions 11 and 12.
For the tail $B_{n,M_{n}}$, note that
$\big|\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}(Y)-\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(Y)\big|\leq\big|\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}(Y)\big|+\big|\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(Y)\big|,$
and that $\sup_{\eta,y}\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}(y)\leq C_{0}+\log(1/h_{n})$ for all large $n$ by Assumptions 10(iv) and 11, while $\sup_{\eta}|\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(Y)|\leq H^{(\ell)}(Y)$ by Assumption 10(vi). Using $\sum_{i}\widehat{w}_{i}^{(\ell)}\leq C_{\ell}$,
$\sup_{\eta}|B_{n,M_{n}}(\eta)|\lesssim\big(C_{0}+\log(1/h_{n})\big)\,P_{n}^{(\ell)}(|Y|>M_{n})+P_{n}^{(\ell)}\big(H^{(\ell)}(Y)\mathbf{1}_{\{|Y|>M_{n}\}}\big).$
By Assumption 13(i) applied to indicators and $H^{(\ell)}$, and Assumption 12,
$\big(C_{0}+\log(1/h_{n})\big)\,P_{n}^{(\ell)}(|Y|>M_{n})=o_{p}(1),\qquad P_{n}^{(\ell)}\big(H^{(\ell)}(Y)\mathbf{1}_{\{|Y|>M_{n}\}}\big)\to_{p}0.$
Thus $\sup_{\eta}|B_{n,M_{n}}(\eta)|=o_{p}(1)$.
Control of $R_{2,n}$. By Assumptions 10(vi) and 13(i), for each fixed $\eta$,
$R_{2,n}(\eta)=P_{n}^{(\ell)}\big(\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}\big)-\mathbb{E}_{f^{(\ell)}}\big[\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(Y)\big]\to_{p}0.$
Continuity of $\eta\mapsto\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(y)$, compactness of $\mathcal{C}\times\Pi_{\min}$, and the envelope $H^{(\ell)}$ imply, by a finite $\varepsilon$–net argument, that $\sup_{\eta}|R_{2,n}(\eta)|\to_{p}0.$
Combining these bounds and taking the supremum over $\eta$ yields the claim. ∎
Proof of Theorem 4 (stage-$\ell$ consistency)
Proof.
By Proposition 3, $\sup_{\eta\in\mathcal{C}\times\Pi_{\min}}|\tilde{\mathcal{L}}_{n}^{(\ell)}(\eta)-\mathcal{L}^{(\ell)}(\eta)|\to_{p}0.$ By Assumption 10(v), $L^{(\ell)}$ is continuous on the compact set $\mathcal{C}\times\Pi_{\min}$ with a unique maximiser $\eta_{\ell}^{*}$. Wald’s argmax theorem yields $\hat{\eta}_{\ell,n}\to_{p}\eta_{\ell}^{*}$. ∎
CLT proof and supporting lemmas
Assumption 15 (Oracle differentiability and information at stage $\ell$).
For each $y$, $\eta\mapsto\ell^{(\ell)}(y;\eta,f_{\mathrm{bg}}^{(\ell),*})$ is twice continuously differentiable on $\mathcal{C}\times\Pi_{\min}$ and differentiation under the integral sign is valid. Define the oracle score and Hessian
$S_{n}^{(\ell)}(\eta):=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}\nabla_{\eta}\ell^{(\ell)}\big(Y_{i};\eta,f_{\mathrm{bg}}^{(\ell),*}\big),\quad H_{n}^{(\ell)}(\eta):=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}\nabla_{\eta}^{2}\ell^{(\ell)}\big(Y_{i};\eta,f_{\mathrm{bg}}^{(\ell),*}\big).$
There exist envelopes $H_{1}^{(\ell)},H_{2}^{(\ell)}\in L^{2}(f^{(\ell)})$ with
$\|\nabla_{\eta}\ell^{(\ell)}(Y;\eta,f_{\mathrm{bg}}^{(\ell),*})\|\leq H_{1}^{(\ell)}(Y),\quad\|\nabla_{\eta}^{2}\ell^{(\ell)}(Y;\eta,f_{\mathrm{bg}}^{(\ell),*})\|\leq H_{2}^{(\ell)}(Y)$
for all $\eta\in\mathcal{C}\times\Pi_{\min}$, and
$\sup_{\eta\in\mathcal{C}\times\Pi_{\min}}\big\|H_{n}^{(\ell)}(\eta)-\mathbb{E}_{f^{(\ell)}}[\nabla_{\eta}^{2}\ell^{(\ell)}(Y;\eta,f_{\mathrm{bg}}^{(\ell),*})]\big\|\ \xrightarrow{p}\ 0.$
Moreover, the information matrix
$J^{(\ell)}(\eta_{\ell}^{*}):=\mathbb{E}_{f^{(\ell)}}\big[\nabla_{\eta}\ell^{(\ell)}(Y;\eta_{\ell}^{*})\nabla_{\eta}\ell^{(\ell)}(Y;\eta_{\ell}^{*})^{\top}\big]=-\,\mathbb{E}_{f^{(\ell)}}\big[\nabla_{\eta}^{2}\ell^{(\ell)}(Y;\eta_{\ell}^{*})\big]$
exists, is finite, and is positive definite.
Assumption 16 (Oracle score CLT at stage $\ell$).
Under the weights $\widehat{w}_{i}^{(\ell)}$ and $f^{(\ell)}$, the oracle score satisfies
$\sqrt{n}\,S_{n}^{(\ell)}(\eta_{\ell}^{*})\ \rightsquigarrow\ \mathcal{N}\big(0,J^{(\ell)}(\eta_{\ell}^{*})\big).$
Assumption 17 (Oracle–equivalence of plug–in score/Hessian).
Let
$\tilde{S}_{n}^{(\ell)}(\eta):=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}\nabla_{\eta}\ell^{(\ell)}\big(Y_{i};\eta,\tilde{f}_{\ell,n}^{\mathrm{FFT}}\big),\quad\tilde{H}_{n}^{(\ell)}(\eta):=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}\nabla_{\eta}^{2}\ell^{(\ell)}\big(Y_{i};\eta,\tilde{f}_{\ell,n}^{\mathrm{FFT}}\big).$
Assume
$\sqrt{n}\,\big\|\tilde{S}_{n}^{(\ell)}(\eta_{\ell}^{*})-S_{n}^{(\ell)}(\eta_{\ell}^{*})\big\|\ \xrightarrow{p}\ 0,\qquad\sup_{\eta\in\mathcal{C}\times\Pi_{\min}}\big\|\tilde{H}_{n}^{(\ell)}(\eta)-H_{n}^{(\ell)}(\eta)\big\|\ \xrightarrow{p}\ 0.$
Proof of Theorem 5 (stage-$\ell$ CLT)
By Theorem 4, $\hat{\eta}_{\ell,n}\to_{p}\eta_{\ell}^{*}$. The first–order condition for the maximiser is $\tilde{S}_{n}^{(\ell)}(\hat{\eta}_{\ell,n})=0$. Let $\tilde{\eta}_{\ell,n}$ lie on the line segment between $\hat{\eta}_{\ell,n}$ and $\eta_{\ell}^{*}$. A Taylor expansion of $\tilde{S}_{n}^{(\ell)}$ around $\eta_{\ell}^{*}$ gives
$0=\tilde{S}_{n}^{(\ell)}(\eta_{\ell}^{*})+\tilde{H}_{n}^{(\ell)}(\tilde{\eta}_{\ell,n})(\hat{\eta}_{\ell,n}-\eta_{\ell}^{*}),$
so
$\sqrt{n}(\hat{\eta}_{\ell,n}-\eta_{\ell}^{*})=-\,\tilde{H}_{n}^{(\ell)}(\tilde{\eta}_{\ell,n})^{-1}\sqrt{n}\,\tilde{S}_{n}^{(\ell)}(\eta_{\ell}^{*}).$
Decompose
$\sqrt{n}\,\tilde{S}_{n}^{(\ell)}(\eta_{\ell}^{*})=\sqrt{n}\,S_{n}^{(\ell)}(\eta_{\ell}^{*})+\sqrt{n}\,\big(\tilde{S}_{n}^{(\ell)}(\eta_{\ell}^{*})-S_{n}^{(\ell)}(\eta_{\ell}^{*})\big).$
By Assumption 16, $\sqrt{n}\,S_{n}^{(\ell)}(\eta_{\ell}^{*})\rightsquigarrow\mathcal{N}(0,J^{(\ell)}(\eta_{\ell}^{*}))$. By Assumption 17, $\sqrt{n}(\tilde{S}_{n}^{(\ell)}(\eta_{\ell}^{*})-S_{n}^{(\ell)}(\eta_{\ell}^{*}))\to_{p}0$ and $\sup_{\eta}\|\tilde{H}_{n}^{(\ell)}(\eta)-H_{n}^{(\ell)}(\eta)\|\to_{p}0.$ Assumption 15 implies $H_{n}^{(\ell)}(\eta)\to_{p}-J^{(\ell)}(\eta)$ uniformly on $\mathcal{C}\times\Pi_{\min}$, so $\tilde{H}_{n}^{(\ell)}(\tilde{\eta}_{\ell,n})\to_{p}-J^{(\ell)}(\eta_{\ell}^{*})$. By Slutsky’s theorem,
$\sqrt{n}(\hat{\eta}_{\ell,n}-\eta_{\ell}^{*})\ \rightsquigarrow\ \mathcal{N}\big(0,J^{(\ell)}(\eta_{\ell}^{*})^{-1}\big).$